Richard has 22 coins with a total value of $4. If the coins are all quarters and dimes, how many of each type of coin does he have?
Answer by ankor@dixie-net.com(22740)

Richard has 22 coins
d + q = 22
d = (22 - q)
:
with a total value of $4.
.10d + .25q = 4
:
If the coins are all quarters and dimes, how many of each type of coin does he have?
:
Substitute (22-q) for d in the above equation:
.10(22-q) + .25q = 4
2.2 - .10a + .25q = 4
-.10q + .25q = 4 - 2.2
.15q = 1.8
q = 1.8/.15
q = 12 quarters
:
That would give us: 22 - 12 = 10 dimes
:
:
Check solutions
.10(10) + .25(12) =
1.00 + 3.00 = 4.00